A body of mass 2 kg is moving on a smooth horizontal surface under the action of a horizontal force F=55+t² (N). Calculate the velocity of the body at t=5s, assuming the body was at rest at t=0. (Answer is 158.3m/s)


1
Expert's answer
2021-04-29T07:29:11-0400

To be given in question

Mass =2kg

Force(F)="55+t^2(N)"

t=5sec

To be asked in question

Velocity of body(V)=?

We know that

"F=ma=m\frac{dv}{dt}\rightarrow(1)"

Put value

"55+t^2=2\frac{dv}{dt}"

"\frac{dv}{dt}=\frac{1}{2}[55+t^2]\rightarrow(2)"


eqution (2)take integration with respect to time

"v=\frac{1}{2}[55t+\frac{t^3}{3}]\rightarrow(3)"


Put(t=5sec)in eqution (3)

"v=\frac{1}{2}[55\times5+\frac{125}{3}]"

"v=\frac{1}{2}[275+\frac{125}{3}]"

"v=\frac{950}{6}=158.33meter/sec"

"v=158.33meter\/sec"